What potential differences is needed to stop an electron that has an initial velocity v=5.0x10^5 m/s ?
1
Expert's answer
2020-03-25T11:00:24-0400

The change of kinetic energy of an electron is equal to work done by an electric field


"\frac{mv^2}{2}=eV."

Hence,


"V=\frac{mv^2}{2e}=\frac{9.1\times 10^{-31}\times (5.0\times 10^5)^2}{2\times 1.6\times 10^{-19}}=0.71 \:\rm V."